OpenStudy (anonymous):

Replace PQ, QR, and PR with the given information. PQ + QR = PR PQ = 2x - 1 QR = 3x + 5 PR = 34 (2x - 1) + (3x + 5) = 34 i know its simple but i don't understand what its asking. i have online school and the computer doesn't explain well

OpenStudy (anonymous):

just solve the equation,find x,then find PQ & QR

OpenStudy (anonymous):

Thats what i'm having problems with the PQ and QR, I might just have to call up my teacher and ask.

OpenStudy (anonymous):

what do you get for x ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

good then PQ=2x-1=2(6)-1=.... QR=3x+5=3(6)+5=....

OpenStudy (anonymous):

so 11+23=34

OpenStudy (anonymous):

yes,but i admit that the question isn't clearly stated

OpenStudy (anonymous):

It' s online school it's expected.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thank you(: